OpenStudy (anonymous):

Rationalize the denominator of 9 over quantity of 3 plus 2 i

OpenStudy (anonymous):

multiply top and bottom by conjugate.

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

Do you know what the conjugate of \[ 3+2i \]is?

OpenStudy (anonymous):

@wio 3-2i

OpenStudy (anonymous):

Okay then do this: \[ \frac{9}{3+2i}\times \frac{3-2i}{3-2i} \]

OpenStudy (anonymous):



OpenStudy (anonymous):

- 18i

OpenStudy (anonymous):